# High-pressure characterization of Ag3AuTe2: Implications for strain-induced band tuning

### v1

@@ III.4 Optical Properties
Figure 7: Tauc plot of Ag3AuTe2 obtained from the diffuse reflectance measurement. The optical band gap energy was estimated by the intersection of the baseline (blue) and fitted line (red) since the baseline was not flat to zero. The approximate band gap is around 0.58 eV.
@@
The above relations with $\gamma=1/2$ results in a Tauc plot with a gentle, rather than a steep rise at approximately 0.6 eV, as shown in Figure 7, consistent with the predicted mBJ ambient pressure indirect gap. However, the PBE gap falls short of the optical band gap by orders of magnitude, as expected due to the tendency of PBE functional to underestimate band gaps. The Arrhenius activation energy $E_{a}$ (0.2-0.3 eV) is lower but comparable to the optical gap. Discrepancies in the experimental gap values may stem from factors such as errors in sample dimensions for transport measurement and a rough approximation of the optical absorption onset.
The correspondence between the optical and mbJ gap suggests promising avenues for exploiting pressure-induced band modulation of Ag3AuTe2 for possible gap size reduction. This merits further investigation of the pressure-induced trend in band gap through in-situ high-pressure UV-Vis measurements.
@@ IV Conclusions
We investigated the synthesis, electronic and optical properties, as well as the compressibility of polycrystalline Ag3AuTe2. High-pressure XRD measurements revealed a pressure-induced structural transition of petzite around 16 GPa, shifting from a cubic to an orthorhombic Ag2Te-type structure. Transport results corroborated Young et al.’s observations,Young et al. (2000) highlighting the semiconductor nature of the telluride with an Arrhenius activation energy of 0.2-0.3 eV, which falls within the predicted gap range by PBE and mBJ DFT calculations. The absorption band edge determined by diffuse reflectance measurements was in excellent agreement with the mBJ gap. First-principles calculations using PBE and mBJ functionals predicted a band gap reduction with decreasing lattice parameters, reaching zero and 0.3 eV gap at -4% strain, respectively. Our high-pressure XRD experiments demonstrated a reduction of over 4% in lattice parameters in Ag3AuTe2 over the measured range of pressure. These findings suggest a viable band engineering avenue for Ag3AuTe2 towards zero gap under applied pressure, encouraging future investigation through $in$ $situ$ high-pressure UV-Vis-IR spectroscopy.

### v2

@@ III.4 Optical Properties
Figure 7: Tauc plot of Ag3AuTe2 was obtained from the diffuse reflectance measurement. The optical band gap energy was estimated by the intersection of the baseline (blue) and fitted line (red) since the baseline was not flat to zero. The approximate band gap is 0.58 eV.
@@
Using the above relations with $\gamma=1/2$ resulted in a Tauc plot with a gentle, rather than a steep rise at approximately 0.6 eV, as shown in Figure 7, consistent with the predicted mBJ ambient pressure indirect gap. However, the PBE gap falls short of the optical band gap by orders of magnitude, as expected due to the tendency of PBE functional to underestimate band gaps. The Arrhenius activation energy $E_{a}$ (0.2 eV) is lower than but comparable to the optical gap in magnitude. Discrepancies in the experimental gap values may arise from factors such as errors in sample dimensions for transport measurement and a rough approximation of the optical absorption onset. The correspondence between the optical and mbJ gap suggests promising avenues for exploiting pressure-induced band modulation of Ag3AuTe2.
@@ IV Conclusions
We investigated the synthesis, electronic and optical properties, as well as the compressibility of polycrystalline Ag3AuTe2. High-pressure XRD measurements revealed a pressure-induced structural transition of petzite around 16 GPa, shifting from a cubic to an orthorhombic Ag2Te-type structure. Transport results corroborated Young et al.’s observations,Young et al. (2000) highlighting the semiconductor nature of the telluride with an Arrhenius activation energy of 0.2 eV, which falls within the predicted gap range by PBE and mBJ DFT calculations. The absorption band edge determined by diffuse reflectance measurements was in excellent agreement with the mBJ gap. First-principles calculations using PBE and mBJ functionals predicted a gap reduction with decreasing lattice parameters, reaching zero and 0.3 eV gap at -4% strain, respectively. Our high-pressure XRD experiments demonstrated a lattice parameter reduction of over 4% in Ag3AuTe2, suggesting a viable band engineering avenue for Ag3AuTe2 towards zero gap under applied pressure.